OpenStudy (anonymous):

Which of these is equivalent to thirteen-eighteenths ? Answer eighteen thirty-firsts thirty-nine sixths twenty-six thirty-firsts twenty-six thirty-sixths

OpenStudy (lgbasallote):

\[\frac{13}{18} \times \frac 22 \implies \frac{13 \times 2}{18 \times 2} \implies \frac{26}{36}\] does that help?

OpenStudy (anonymous):

yassss alot